OpenStudy (anonymous):

jason bought a shirt that was 40% off the original price. which fraction is equivalent to 40%? a. 2/5 b. 40/90 c. 1/40 d. 4/25

OpenStudy (anonymous):

1/10= 10% thus 4/10=40% simplifying =2/5